Circular No. 8158

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003gg IN IC 1321 Further to IAUC 8156, M. Moore, M. Papenkova, and W. Li report the

LOTOSS discovery of an apparent supernova (mag ∼ 18.5) on unfiltered KAIT images on June 27.4 and 28.4 UT. SN 2003gg is located at α = 20^{h}28^{m}10^{s}.68, δ = −18^{o}17′32′′.4 (equinox 2000.0), which is 5′′.8 west and 2′′.9 south of the nucleus of IC 1321. A KAIT image taken on 2002 July 9.4 showed nothing at this position (limiting mag ∼ 19.0). A. V. Filippenko and R. Chornock, University of California, Berkeley, report that inspection of CCD spectra (range 320–900 nm) obtained on June 29 with the Keck-I telescope (+ LRIS) reveals that SN 2003gg is of type II, probably 1–2 months after the explosion. Hydrogen Balmer lines, Na D, and Fe II lines exhibit well-developed P-Cyg profiles.

NO NOVA IN NGC 6822 Filippenko and Chornock report that inspection of CCD spectra (range

320–990 nm), obtained as above on June 29 UT, reveals that the apparent nova in NGC 6822 (cf. IAUC 8141) is actually a very late M-type star in that galaxy. Strong TiO bands are present, as is narrow Hα emission. Such stars are known to sometimes be highly variable, such as Mira variables.

VARIABLE STAR NEAR UGC 10700 Filippenko and Chornock also note that CCD spectra show that the

suspected variable star near UGC 10700 (cf. IAUC 8153) is indeed a Galactic cataclysmic variable. Strong, narrow, double-peaked (separation ∼ 500 km/s) hydrogen Balmer emission lines are present. He I and the Ca II near-infrared triplet emission are visible, as well.

SUPERNOVAE 2003ds, 2003ev, AND 2003fd Filippenko and Chornock add that inspection of CCD spectra, ob-

tained as above on June 29 UT with the Keck-I telescope, reveals that SN 2003fd (IAUC 8147) is of type Ia, perhaps 2 weeks past maximum brightness. SN 2003ev (IAUC 8140) is indeed of type Ic, as suspected by Hamuy and Phillips (IAUC 8142). SN 2003ds (IAUC 8120, 8129) is also of type Ic.

SUPERNOVA 2003gd IN M74 Prediscovery unfiltered-CCD magnitude reported by A. Klotz, Guital-

ens, France: June 11.12 UT, V = 13.65 ± 0.3.

2003 June 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green